OpenStudy (anonymous):

Express 17 1/2% as decimal?

OpenStudy (anonymous):

Remember that 17.5% = 17.5/100 I'll let you figure the rest out.

OpenStudy (anonymous):

how's the working for this?

OpenStudy (anonymous):

@Quaternions ?

OpenStudy (anonymous):

I'm sorry. What? I don't understand the question. Do you mean how to divide 17.5 by 100?

OpenStudy (anonymous):

A percent X% can always be represented as a fraction of X/100. And if you need to have a decimal approximation, then divide by 100 (or move the decimal place two places to the right. For example 40% = 40/100 = 0.40

OpenStudy (anonymous):

thankyou